# Out-of-Distribution Detection using Neural Activation Prior

### 01 Mar 2024 v2

@@ 4.1 Basics
First, we will provide a brief overview of typical settings for out-of-distribution (OOD) detection in image classification networks. Typically, classification networks use in-distribution (ID) data, that is, known training data sets, to train a classification model. Once training is completed, the model’s parameters are fixed, enabling it to effectively classify the categories within the training data set. In the testing phase, to identify OOD samples, researchers usually introduce a scoring function into the model. During the inference, samples mixed with OOD data are fed into this trained model. The model not only classifies each sample, but also uses a scoring function to generate a score for each one. This score is used to evaluate whether a sample belongs to an in-distribution class in the training set, or an unknown out-of-distribution class.
@@ 4.2 Design of scoring function
where $C$ represents the number of channels before global pooling. In this formula, a small constant $\epsilon>0$ is added to ensure the numerical stability of the computation.
@@ 5 Experiments
Table 1: Comparison with competitive post hoc out-of-distribution detection method on CIFAR [19] benchmarks. All values are percentages and are averaged over 6 OOD test datasets. Except for our proposed method, all other results are taken from [8]. The best result is highlighted in bold. Methods marked with an asterisk (*) are employed with the Energy Score [23].
Method CIFAR-10 CIFAR-100
FPR95 AUROC FPR95 AUROC
MSP [11] 48.73 92.46 80.13 74.36
ODIN [22] 24.57 93.71 58.14 84.49
Mahalanobis [21] 31.42 89.15 55.37 82.73
Energy [23] 26.55 94.57 68.45 81.19
ReAct* [34] 26.45 94.67 62.27 84.47
DICE* [33] 20.83 95.24 49.72 87.23
ASH-B* [8] 20.23 96.02 48.73 88.04
ASH-S* [8] 15.05 96.91 41.40 90.02
NAP-E (Ours) 9.02 98.15 32.61 92.84
NAP-A (Ours) 11.14 97.48 35.40 91.21
NAP-R (Ours) 9.18 98.02 25.71 93.18
Table 2: OOD detection results on ImageNet-1k [6]. All values are percentages. Except for our proposed method, all other results are taken from [8]. The best result is highlighted in bold. Methods marked with an asterisk (*) are employed with the Energy Score [23].
@@
iNaturalist [38] SUN [39] Places [45] Textures [5]
@@
NAP-E (Ours) 29.90 94.47 39.69 90.46 55.17 85.15 11.74 97.28 34.12 91.84
NAP-A (Ours) 28.34 94.69 36.29 91.97 51.25 87.27 12.36 97.10 32.06 92.76
NAP-R (Ours) 24.58 95.55 38.47 91.12 53.32 86.24 9.57 97.60 31.49 92.63
5.1 Evaluation on CIFAR benchmarks
Implementation Details. In our experiments, consistent with recent studies [34, 33, 8], we utilized 10,000 test images from both CIFAR-10 [19] and CIFAR-100 [19] as in-distribution data. To gauge the performance of the model, six widely-used OOD datasets were employed as benchmarks. These datasets include SVHN [27], Textures [5], iSUN [40], LSUN-Crop [42], LSUN-Resize [42], and Places365 [45]. As for pre-trained model, we employed DenseNet [13], and we follow the training setting of DenseNet introduced in [33].
[…]
@@ 5.2 Evaluation on ImageNet
Experimental Results. Table 2 shows the comparison results of NAP with other post hoc OOD detection methods on the ImageNet-1k [6] benchmark. As shown in the table, our method outperforms all other methods on average on the ImageNet-1k dataset [6], achieving state-of-the-art performance. We find that our method is less effective on the Places dataset [45]. We argue that this is because the scene in Places [45] is complex and contains some patterns that can cause large response values. These patterns are close to the patterns that the original network focuses on, so it is relatively difficult to distinguish them.
@@
Figure 4: Investigating the trade-offs between In-Distribution (ID) classification accuracy and Out-Of-Distribution (OOD) detection rate on CIFAR benchmarks [19] across various methods. All methods and experiments are implemented by us. Methods prefixed with ’NAP’ are visually distinguished, highlighted in various shades of red in the figures.
We also note that all three of our implementations outperform the state-of-the-art methods on the iNaturalist [38] and Textures [5] datasets. Intuitively, most of the pictures in the Textures [5] dataset are of a texture nature, so there is a small probability of obtaining a large response value. And while samples in iNaturalist have a relatively simple background, and the animals and plants in the foreground have certain semantic differences from the samples in ImageNet-1k, therefor, a sample in iNaturalist [38] is not likely to trigger large response values. The reason why existing methods do not perform well on this dataset may be due to their focus on activation values after global pooling. We speculate that although texture pictures cannot cause large activation values, they can excite small-amplitude noise on the entire feature map. Since large response values tend to only appear in a small area on the feature map, after global pooling, in-distribution samples and texture class samples are likely to obtain similar average activation values. This may compromise the separability of the two types of samples, causing existing work to perform poorly on such samples.
5.3 Evaluation on Pareto Frontier of ID accuracy and OOD detection
The goal of OOD detection methods is to provide the best OOD detection performance without compromising the inherent classification performance of the original model. While some existing methods improve OOD detection performance, they also impair the model’s classification ability within the distribution. Although the authors of these methods claimed in their papers that the damage was minor, when we tried to combine different methods, the subtle changes to the model by different methods eventually led to a significant decrease in the classification ability of the original model. This cumulative damage hinders the possibility of combining multiple methods for OOD detection. Based on NAP, our proposed method preserves the model’s classification ability and is simple to implement, thereby offering a cost-free enhancement in OOD detection performance. Figure 3(a) and Figure 3(b) highlight the trade-off between in-distribution (ID) classification accuracy and out-of-distribution (OOD) detection rate (AUROC) on the CIFAR benchmark [19]. Our proposed NAP family demonstrates the optimal balance of this trade-off, demonstrating superior performance along the Pareto front.
[…]
Table 3: Comparison with competitive out-of-distribution score method on CIFAR benchmarks [19] and ImageNet-1k [6] dataset. CIFAR results are averaged across 6 different OOD tasks and ImageNet-1k results are averaged across 4 different OOD tasks.
@@ 5.3 Evaluation on Pareto Frontier of ID accuracy and OOD detection
(a) After first convolution layer
(b) In Transaction Block 1
(c) In Transaction Block 2
(d) Before Global Pooling
Figure 5: Activation distribution at different positions within the DenseNet architecture [13] applied to CIFAR-10 [19] and Places365 datasets [45]. For this analysis, four specific locations within the network were chosen: (a) after the first convolution layer, (b) just before the pooling operation in the first transition block, (c) just before the pooling operation in the second transition block, and (d) right before the final global pooling layer. It is observed that with increasing depth, the separability between in-distribution and out-of-distribution samples becomes more pronounced in the two-dimensional space defined by the maximum and average activation values.
@@ 5.5 Ablation studies
5.5.1 Ablation on different parameter $w$.
Table 4: Ablation results on different $w$ of NAP-R on CIFAR [19] benchmarks and ImageNet-1k [6] dataset. CIFAR results are averaged across 6 different OOD tasks and ImageNet-1k results are averaged across 4 different OOD tasks.
$w$ CIFAR-10 CIFAR-100 ImageNet-1k
FPR95 AUROC FPR95 AUROC FPR95 AUROC
@@
In Table 4, we show the various OOD detection performances of NAP-R on CIFAR10 [19], CIFAR100 [19] and ImagetNet-1k [6] by varying the threshold $w$. Experimental results show that by choosing an appropriate $w$, we can obtain better OOD detection results by combining NAP with the existing OOD scoring function. This intuitively indicates that our proposed prior is orthogonal to and effectively complements existing methods. See the Appendix for the performance impact of changes in $w$ for other implementations.
5.5.2 Does NAP Work across all network layers?
In our study, we visualized the distribution of activations at different positions in the DenseNet architecture applied to the CIFAR-10 [19] and Places365 datasets [45]. Four specific locations were chosen for this analysis: after the first convolutional layer, before the pooling operation in the first transition block, before pooling in the second transition block, and before the final global pooling layer. We selected some typical cases from the experimental results to display in Figure 5. For a more comprehensive set of visualization results, please see the Appendix. From Figure 5, we can observe that as the position becomes deeper, the separability of in-distribution samples and out-of-distribution samples becomes more and more obvious in the two-dimensional space spanned by the maximum and average activation values. This is because in the first few layers of the network, neurons tend to capture some low-level features, and such features are likely to be present in both ID and OOD samples. Therefore, as seen in Figures 4(a) and 4(b), the two lines intersect. As we examine deeper neurons in the network, the features captured by neurons become features containing certain semantic information. Therefore, as shown in  4(c) and 4(d), data within the distribution with specific semantic information can obtain relatively larger response values. This suggests that our proposed Neural Activation Prior is more effective towards the end of the network, particularly before the global pooling layer.

### 06 Mar 2024 v3

@@ 4.1 Basics
First, we will provide a brief overview of typical settings for OOD detection in image classification networks. Typically, classification networks use ID data, that is, known training data sets, to train a classification model. Once training is completed, the model’s parameters are fixed, enabling it to effectively classify the categories within the training data set. In the testing phase, to identify OOD samples, researchers usually introduce a scoring function into the model. During the inference, samples mixed with OOD data are fed into this trained model. The model not only classifies each sample, but also uses a scoring function to generate a score for each one. This score is used to predict whether a sample belongs to an ID class in the training set, or an unknown OOD class.
@@ 4.2 Design of Scoring Function
where $C$ represents the number of channels before global pooling. Note that a small constant $\epsilon>0$ is added to ensure the numerical stability of the computation.
Scoring Function for Transformers. Consistent with the NAP score used in CNNs, we calculate the mean and maximum values of the attention that the cls token has towards all other tokens. The attention vector, denoted as $A$, has a dimensionality of $(l+1)$, where $l$ represents the sequence length. To maintain consistency with the NAP score calculation method used in CNN networks, we would typically divide the maximum value by the mean. However, we note that the mean value of the attention vector is always $1/(l+1)$, rendering the denominator superfluous. Therefore, for simplicity, we design the NAP score function for Transformers as $S_{NAP}^{Former}=\text{Max}(A)$.
@@ 5 Experiments
5.1 Evaluation on CIFAR Benchmarks
Implementation Details. In our experiments, consistent with recent studies [33, 34, 8], we utilized 10,000 test images from both CIFAR-10 [20] and CIFAR-100 [20] as ID data. To gauge the performance of the model, six widely-used OOD datasets were employed as benchmarks. These datasets include SVHN [28], Textures [5], iSUN [41], LSUN-Crop [43], LSUN-Resize [43], and Places365 [46]. As for pre-trained model, we employed DenseNet [15], and we follow the training setting of DenseNet introduced in [34].
Table 1: Comparison with competitive post-hoc OOD detection method on CIFAR benchmarks. All values are percentages and are averaged over 6 OOD test datasets. Except for our proposed method, all other results are taken from [8]. The best result is highlighted in bold. Methods marked with an asterisk (*) are employed with the Energy Score [24]. Note: A. = Area Under the ROC Curve; F. = False Positive Rate at 95% True Positive Rate. Methods include MSP [11], ODIN [23], Mahalanobis distance [22], Energy [24], ReAct [33], DICE [34], and ASH [8].
Method MSP ODIN Maha. Energy ReAct DICE ASH-B ASH-S NAP-E NAP-A NAP-R
CIFAR-10 F. $\downarrow$ 48.73 24.57 31.42 26.55 26.45 20.83 20.23 15.05 9.02 11.14 9.18
A. $\uparrow$ 92.46 93.71 89.15 94.57 94.67 95.24 96.02 96.91 98.15 97.48 98.02
CIFAR-100 F. $\downarrow$ 80.13 58.14 55.37 68.45 62.27 49.72 48.73 41.40 32.61 35.40 25.71
A. $\uparrow$ 74.36 84.49 82.73 81.19 84.47 87.23 88.04 90.02 92.84 91.21 93.18
[…]
Table 2: OOD detection results on ImageNet-1k [6]. All values are percentages. Except for our proposed method, all other results are taken from [8]. Methods marked with an asterisk (*) are employed with the Energy Score [24].
@@ 5.1 Evaluation on CIFAR Benchmarks
iNat. [38] SUN [40] Places [46] Textures [5]
@@
NAP-E 29.90 94.47 39.69 90.46 55.17 85.15 11.74 97.28 34.12 91.84
NAP-A 26.26 95.10 32.89 92.77 48.69 87.92 11.60 97.32 29.86 93.28
NAP-R 24.58 95.55 38.47 91.12 53.32 86.24 9.57 97.60 31.49 92.63
@@ 5.2 Evaluation on ImageNet
Experimental Results. Table 2 shows the comparison results of NAP with other post-hoc OOD detection methods on the ImageNet-1k [6] benchmark. As shown in the table, our method outperforms all other methods on the ImageNet-1k dataset [6], achieving state-of-the-art performance. We have observed that our method brings significant performance improvements on the iNaturalist and Textures datasets, which is particularly notable. Intuitively, most of the pictures in the Textures [5] dataset are of a texture nature, so there is a small probability of obtaining a large response value. And while samples in iNaturalist have a relatively simple background, and the animals and plants in the foreground have certain semantic differences from the samples in ImageNet-1k, therefor, a sample in iNaturalist [38] is not likely to trigger large response values. The reason why existing methods do not perform well on this dataset may be due to their focus on activation values after global pooling. We speculate that although texture pictures cannot cause large activation values, they can excite small-amplitude noise on the entire feature map. Since large response values tend to only appear in a small area on the feature map, after global pooling, ID samples and texture class samples are likely to obtain similar average activation values. This may compromise the separability of the two types of samples, causing existing work to perform poorly on such samples.
5.3 Evaluation on Transformer
Table 3: OOD detection results on ViT-B/16 using ImageNet-1k as ID data. All values are percentages. Except for our proposed method, all other results are taken from [8]. The best result is highlighted in bold. Methods marked with an asterisk (*) are employed with the Energy Score [24].
Method OOD Datasets Average
iNat. [38] SUN [40] Places [46] Textures [5]
FPR95 AUROC FPR95 AUROC FPR95 AUROC FPR95 AUROC FPR95 AUROC
Energy [24] 64.08 79.24 72.77 70.25 74.30 68.44 58.46 79.30 67.40 74.31
NAP-E 60.97 80.77 64.05 77.34 69.34 73.30 45.04 86.93 59.85 79.58
MSP [11] 51.47 88.16 66.53 80.93 68.65 80.38 60.21 82.99 61.72 83.12
NAP-M 47.09 88.23 59.45 82.78 63.38 80.48 47.70 87.93 54.40 84.85
Following the experimental setup described in Section 5.2, we conduct experiments on the Vision Transformer [9] (ViT-B/16) using ImageNet-1k as the ID dataset. Energy [24] and MSP [11] were selected as baseline methodologies for this analysis. The study further explores the enhancement of these baseline methods through the integration of NAP, resulting in two variants: NAP-E and NAP-M. Comparative results detailed in Table 3 demonstrate that the NAP method substantially boosts the performance on Transformer architectures beyond the baselines, affirming the utility and versatility of NAP within such contexts.
@@ 5.3 Evaluation on Transformer
Figure 4: Investigating the trade-offs between ID classification accuracy and OOD detection AUROC on CIFAR benchmarks [20] across various methods. All methods and experiments are implemented by us. Methods prefixed with ’NAP’ are visually distinguished, highlighted in various shades of red in the figures.
5.4 Evaluation on Pareto Frontier of ID accuracy and OOD Detection Performance
Some existing methods negatively affect ID accuracy; however, we have found that integrating these methods with the NAP approach can mitigate or reduce this impact, achieving a more optimal balance. The NAP approach establishes an ideal equilibrium between ID classification accuracy and OOD detection efficacy (measured in AUROC), thereby positioning it favorably on the Pareto front for superior performance. This demonstrates our method’s ability to enhance OOD detection without additional costs while maintaining the model’s classification performance, as illustrated in Figures 3(a) and  3(b) using the CIFAR benchmark [20].
Table 4: Comparison with other single OOD score on CIFAR benchmarks [20] and ImageNet-1k [6] dataset. CIFAR results are averaged across 6 different OOD tasks and ImageNet-1k results are averaged across 4 different OOD tasks.
@@ 5.5 Ablation Studies
[…]
Table 5: Ablation results on different $w$ of NAP-E on CIFAR [20] benchmarks and ImageNet-1k [6] dataset.
Table 6: Ablation results on different $w$ of NAP-R on CIFAR [20] benchmarks and ImageNet-1k [6] dataset.
$w$ CIFAR-10 CIFAR-100 ImageNet
FPR95 AUC FPR95 AUC FPR95 AUC
0.0 26.57 92.45 54.91 85.86 49.91 83.07
0.1 17.94 95.44 47.95 88.85 47.35 85.01
0.2 11.64 97.19 40.24 91.14 44.39 86.99
0.3 9.26 97.96 34.24 92.50 41.18 88.84
0.4 9.02 98.15 32.61 92.84 37.99 90.37
0.5 9.71 98.00 35.13 92.19 35.25 91.41
0.6 11.85 97.63 42.30 90.63 34.12 91.84
0.7 14.56 97.08 49.50 88.48 35.50 91.61
0.8 19.20 96.37 57.11 86.05 40.01 90.69
0.9 22.31 95.55 63.21 83.59 48.05 89.02
1.0 26.58 94.63 68.29 81.23 58.87 86.60
$w$ CIFAR-10 CIFAR-100 ImageNet
FPR95 AUC FPR95 AUC FPR95 AUC
@@
Table 6: Ablation results on different $w$ of NAP-R on CIFAR [20] benchmarks and ImageNet-1k [6] dataset.
Ablation on Different Parameter $w$. In Table 6 and 6, we show the various OOD detection performances of NAP-E and NAP-R on various datasets by varying the threshold $w$. It is important to note that our approach is essentially an ensemble approach. Generally, in ensemble approaches, it is advisable to avoid setting $w$ extremely close to 0 or 1, to ensure effective contribution from all constituent models especially when the models are carefully chosen, diverse and useful. As an ensemble method, we provide the results of $w$ over the entire interval [0,1] in Table 6 to show that our ensemble can always achieve significant improvements over the existing baseline and NAP. Note that these results are not provdied to not to demonstrate the sensitivity of NAP to $w$. Furthermore, Table 6 show that near-optimal (if not optimal) performance is attainable around the mid-range values, especially near 0.5. As $w$ is the only hyper parameter involved in our method, it allows us to employ a binary search to to efficiently determine a desired value for $w$. The significant improvements in ensemble results verify that our method is orthogonal to existing energy-based approaches, which is one of the main contributions of this paper to the community. See the Appendix for more details about the performance impact of changes in $w$.
(a) After conv layer 1
(b) In Trans Block 1
(c) In Trans Block 2
(d) Before Pooling
Figure 5: Activation distribution at different positions within the DenseNet architecture [15] applied to CIFAR-10 [20] and Places365 datasets [46]. For this analysis, four specific locations within the network were chosen: (a) after the first convolution layer, (b) just before the pooling operation in the first transition block, (c) just before the pooling operation in the second transition block, and (d) right before the final global pooling layer. It is observed that with increasing depth, the separability between ID and OOD samples becomes more pronounced in the two-dimensional space defined by the maximum and average activation values.
Does NAP Work Across All Network Layers? To assess the efficacy of the NAP across various layers within the DenseNet architecture, we conduct a detailed examination of activation distributions utilizing the CIFAR-10 [20] and Places365 datasets [46]. Analysis focused on four pivotal points within the network: post-first convolutional layer, pre-pooling in the first and second transition blocks, and immediately preceding the final global pooling layer. Selected instances from the empirical findings are illustrated in Selected examples from our tests are shown in Figure 5, with more detailed visuals in the Appendix. From these figures, it’s clear that the deeper the layer, the easier it is to tell apart ID and OOD samples based on the maximum and average activation values. In the network’s early layers, neurons pick up basic features common to both ID and OOD samples, which explains why the lines cross in Figures 4(a) and 4(b). Moving to deeper layers, neurons start capturing more complex, meaningful features. As seen in Figures 4(c) and 4(d), ID samples with specific meanings trigger larger responses. This indicates that NAP works best closer to the end of the network, especially before the global pooling layer, highlighting its value in making models more reliable for OOD detection.